OpenStudy (anonymous):

The container that holds the water for the football team is 3/10 full. After pouring in 5 gallons of water, it is 1/2 full. How many gallons can the container hold?

OpenStudy (acxbox22):

let x be the amount of water it can hold 3/10x+5=1/2x 3/10x+5=5/10x 5=2/10x 5=1/5x x=?